OpenStudy (anonymous):

Solve for X Y=x-13 Y=16 simplify your answer as much as possible

OpenStudy (anonymous):

i know its the same as last time but u see i dont know wat to do when u subtract it

TheSmartOne (thesmartone):

Y = 16 Plug that in to the first equation which is y=x-13

TheSmartOne (thesmartone):

So basically where you see y in the first equation, just put 16 in its place.

OpenStudy (anonymous):

so like this x-13+16

OpenStudy (anonymous):

ah i suck at math

TheSmartOne (thesmartone):

First just replace y with 16 in the equation y = x - 13

TheSmartOne (thesmartone):

just remove y and put 16 in its place.

OpenStudy (anonymous):

oh 16=x-13

OpenStudy (anonymous):

and then i add 13 to both sides?

TheSmartOne (thesmartone):

Exactly!

TheSmartOne (thesmartone):

So 16 + 13 = ?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

is that my answer?

TheSmartOne (thesmartone):

It is! :D

OpenStudy (anonymous):

thank you so much